OpenStudy (anonymous):

cot(sin^-1(-2/3)) Could someone please show me the steps to solving this without a calculator?

OpenStudy (debbieg):

Ahhhh, this is an interesting problem. Well, do you know how to use the fundamental ID, along with knowledge of the quadrant that an angle is in, to find the cos(x), given the sin(x)?

OpenStudy (debbieg):

That is, if I said that sin(x)=# and x is in quadrant 4, find cos(x) would you know how to do that?

OpenStudy (anonymous):

yup :)

OpenStudy (debbieg):

ok, now look at the "inner" part of the expression above. \[\Large \sin^{-1}(-2/3)\] is an angle x such that: sin(x) = ? and sin(x) is in what quadrant?

OpenStudy (debbieg):

(you know the answer to both, by the definition of the inverse sine function) :)

OpenStudy (debbieg):

So once you answer those 2 questions, you will be able to find cos(x) using the fundamental ID. That means you will have sin(x) and cos(x), from which I bet you know how to find cot(x). :)

OpenStudy (anonymous):

how do you find arcsin (-2/3) without a calculator though?

OpenStudy (debbieg):

That is the beauty of this problem. You don't NEED to find it. Let \(\Large x=\sin^{-1}(-2/3)\) Remember what that inverse function means... it just means "the angle that has that sine value, and is in the range of arcsin"

OpenStudy (debbieg):

So because the range of arcsin is limited, you KNOW the quadrant that x lies in. And you have the sin(x), right? You don't NEED x, you just need to know sin(x) and cos(x) to get cot(x). You are given sin(x), and enough info to find cos(x) with Fund'l ID. :)

OpenStudy (debbieg):

Because with \(\Large x=sin^{-1}(-2/3)\), the problem is simply find \(\Large cot(x)\).

OpenStudy (anonymous):

Wait i think i get it. The answer is \[-\sqrt{5/2}\].

OpenStudy (anonymous):

whoops i mean \[\sqrt{5}/-2\]

OpenStudy (anonymous):

and also thank you very much :)

OpenStudy (debbieg):

Haha... was just about to say that. :) That's what I get too! :)

OpenStudy (debbieg):

You're quite welcome, happy to help. I liked this problem, will have to remember it. :)